OpenStudy (anonymous):

Identify the 42nd term of an arithmetic sequence where a1 = -12 and a27 = 66

OpenStudy (dls):

Solve them simultaneously,you will get two equations Use elimination Bingo :)

OpenStudy (anonymous):

i have no clue what i'm even doing to eliminate :(

OpenStudy (anonymous):

a27 = -12 + 26d = 66

OpenStudy (anonymous):

find d

OpenStudy (anonymous):

You are given with first term = -12 and a(27) = 66 So: Use : \[\large a_n = a + (n-1)d\] So: \[66 = -12 + (27-1)d \implies 66 = -12 +26d\] \[d = \frac{78}{26} = 3\] So now you have to find 42th terms: So n= 42 a = -12 and d = 3 \[\large a_{42} = -12 + (42-1 ) (3)\] Solve this now..

OpenStudy (anonymous):

okay @Yahoo! , lets see .... 66+12=78, and 78/26 = 3?

OpenStudy (anonymous):

yes go ahead..

OpenStudy (anonymous):

ok now d = 3 follow watereyes

OpenStudy (anonymous):

a42 = -12 + 41d

OpenStudy (anonymous):

subs for d -12 + 41*3 =??

OpenStudy (anonymous):

@waterineyes so then its.. 111?

OpenStudy (anonymous):

-12 + 123 = 111

OpenStudy (anonymous):

Let me check : -12 + 41*3 = 123 -12 = 111 Yes you are right.. Well Done..

OpenStudy (anonymous):

so thats ur 42 term

OpenStudy (anonymous):

great ! thanks guys :)

OpenStudy (anonymous):

Yahoo no need..